The Community College basketball team made 51 field goals in a game, some were 2 pointers, and some were 3 pointers. They score 130 points. How many 2 point and 3 point baskets were made?
Answer by macston(5194)

.
x=number of two point goals; y=number of three point goals
.
x+y=51
y=51-x
.
2x+3y=130
2x+3(51-x)=130
2x+153-3x=130
153-x=130
23=x
ANSWER 1: They made twenty three 2 point baskets.
.
y=51-x=51-23=28
ANSWER 2: They made twenty eight 3 point baskets
.
CHECK:
2x+3y=130
2(23)+3(28)=130
46+84=130
130=130